OpenStudy (samsan9):

Please help me solve this problem!!!? find the exact value of tan 22.5 using the half angle identity?

OpenStudy (anonymous):

45

OpenStudy (samsan9):

ok i know that i times it by 2 and then ?

OpenStudy (anonymous):

tanθ=2tan(θ/2)1+tan2(θ/2)

OpenStudy (samsan9):

is that the hanlf angle formula?

OpenStudy (samsan9):

i know i am supposed to get \[\sqrt{2}-3 \]

OpenStudy (samsan9):

but idk how :/

OpenStudy (whpalmer4):

well, 22.5 is half of 45, right?

OpenStudy (samsan9):

yes

OpenStudy (whpalmer4):

the half angle formula is \[\tan^2( u) = \frac{1-\cos(2u)}{1+\cos(2u)}\]so if we set \(u = 22.5^\circ\), \(2u= 45^\circ\) and you can compute the value of \(\cos(45^\circ)\) without any trouble, right?

OpenStudy (samsan9):

is their a square root in this formula?

OpenStudy (whpalmer4):

now, that will give you \(\tan^2 (22.5^\circ)\) and you just take the square root to get \(\tan(22.5^\circ)\)

OpenStudy (whpalmer4):

by the way, \(\tan(22.5^\circ)\) is not \(\sqrt{2} - 3\) but rather \(\sqrt{2}-1\)

OpenStudy (samsan9):

oh ok i wrote that wrong

OpenStudy (samsan9):

thank you :D

OpenStudy (samsan9):

ok i am still a little confused on getting the answer

OpenStudy (whpalmer4):

what do you have so far?

OpenStudy (samsan9):

i have \[\sqrt(1}+\sqrt{2/2}\frac{ 2/2 }{-\sqrt{2/2 ? }\]

OpenStudy (whpalmer4):

do you perhaps mean \[\large \frac{1-\frac{\sqrt{2}}2}{1+\frac{\sqrt{2}}2}\]?

OpenStudy (samsan9):

yes

OpenStudy (whpalmer4):

okay, so rationalize the denominator by multiplying both numerator and denominator by the conjugate of the denominator, which is just the denominator with the sign of the radical flipped

OpenStudy (whpalmer4):

you're setting up a difference of squares...

OpenStudy (samsan9):

\[1-\sqrt{2}\]

OpenStudy (whpalmer4):

no... denominator is \[1+\frac{\sqrt{2}}2\]conjugate is \[1-\frac{\sqrt{2}}2\] \[(a+b)(a-b) = a^2-b^2\]let \[a =1\]\[b=\frac{\sqrt{2}}{2}\]\[(a-b)(a+b) = (1)^2-(\frac{\sqrt{2}}2)^2 = 1 - \frac{2}{4} = \frac{1}{2}\] so the denominator of the rationalized fraction is 1/2

OpenStudy (whpalmer4):

numerator is \[(1-\frac{\sqrt{2}}2)^2\]giving us \[\tan^2(22.5^\circ) = \frac{(1-\frac{\sqrt{2}}{2})^2}{\frac{1}{2}}=\]

OpenStudy (whpalmer4):

and then you take the square root of that \[\sqrt{\tan^2(22.5^\circ) }= \sqrt{\frac{(1-\frac{\sqrt{2}}{2})^2}{\frac{1}{2}}} =\frac{1-\frac{\sqrt{2}}2}{\sqrt{\frac{1}{2}}} = \sqrt{2}(1-\frac{\sqrt{2}}{2})\]

OpenStudy (samsan9):

alright thanks i think i understood it aliitle bit more goodnight